SunWindow V2.0

A general workspace managment and enhancment tool Edition 2, December 1993

Bernhard Scholz

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993 Bernhard Scholz.

This is the second edition of the SunWindow manual.

Published by Bernhard Scholz Steinbruchstr. 9, 82398 Polling Germany

Printed copies are NOT available, (yet).

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by Bernhard Scholz.

Warning: The information described in this manual may contain errors or bugs, and may not function as described. All information is subject to enhancement or upgrade for any reason including to fix bugs, add features, or change performance. As with all software upgrades, full compatibility, although a goal, cannot be guaranteed, and is in fact unlikely.

Amiga is a registered trademark of Commodore-Amiga, Inc. Amiga 1000, AmigaDOS, Amiga Workbench, Amiga Kickstart, AmigaGuide are trademarks of Commodore. Commodore is a registered trademark of Commodore Electronics Limited.

SunWindow V2.0

(Revision 026)

28 December 1993

dedicted to Antje 2

SunWindow Copying Conditions

SunWindow is copyrighted by Bernhard Scholz. The program currently distributed includes portions of the ReqTools package \begin{matrix}c \\ ◯ \end{matrix} by Nico Francois. For more informations on the ReqTools package, please read the ‘REQTOOLS.readme’ file on the distribution disk.

SunWindow is distributed under the concepts of SHAREWARE. This means: If you are using SunWindow frequently (I understand a period of time, not longer than four weeks), you have to be registered as a user of SunWindow by Bernhard Scholz.

Permission is granted to make verbatim copies of SunWindow of the whole package, and process it with any packer.

Permission is granted to distribute SunWindow in a non commercial (this includes other SHARE-WARE products) way provided the copyright notice and this permission notice are preserved in all copies under the conditions of verbatim copying.

Permission is granted to distribute SunWindow on public electronic networks, on public mailboxes or disk based PD-series, without profit for the distributor, and the conditions of verbatim copying providing the copyright notice and this permission notice are preserved.

1 Introduction To SunWindow

In this chapter the concept of SunWindow is described in brief. You find informations about whether you can use the program or not, what are the benefits and other useful things, you might want to know, to get a fast impression of what SunWindow is capable to do.

1.1 The Concept Of SunWindow

The user interface the Amiga offers to an user is very nice and was something revolutionary when the Amiga started its career. The Workbench is the main working area, and Intuition — a part of the operating system — serves for the graphical-user-interface (GUI).

Many years have gone now, since the first release of the Amiga’s GUI, but only since release V2.0 of the OS some great enhancements to the GUI were added, but most of them refer to the lookout of the graphics. Also new features were added: the public screens (screens, where any programm can open its windows) and autoscrolling screens (screens, which are bigger than the visible size will scroll with the mouse to make other parts of such a big screen visible). Although these two features are really great, they are not used very often. To make the use a little bit easier and intuitive to the user, and nearby enhance the control over the GUI, SunWindow was born.

SunWindow offers several things to make the life easier with your daily work. Some of these features are now described:

• SunWindow offers several hotkeys for controlling windows and screens, therefore it serves as a kind of replacement for FKey (by Commodore). Windows can be moved, sized, zipped and closed by hotkeys. Sizing has been enhanced in a way which makes it now possible to size a window in any direction. Screens can be moved and shuffled, as you would do it with the mouse.

• Public screens are automatically under SunWindow’s control (and they are the only kind of screens on which all of SunWindow’s screen and window operations will take affect). If such a screen is bigger than the displayable area of your monitor, SunWindow offers you the possibility to display the whole screen as a thumbnail display in a seperate small window. On this display all of the screen’s windows are redisplayed, and you can perform several actions on these windows, like bringing a window to front, making a selected window visible or just moving the windows around to get the screen cleaned up.

• SunWindow offers you a SunMouse feature. This means, if your mouse stops over a part of a window, the window will become the active window, could be made visible or just brought to front.

• SunWindow offers you also the unique possibility to remember up to ten different positions of your mouse (also on different screens). With just a fingertip on a hotkey you can jump back to any or these remembered positions, without looking through different screens and searching around a big screen to find the desired window/position.

• On big public screens new windows may not be visible, when they are opened. This can be controlled by SunWindow, so that the user will notice every new window, by just centering the display around every new window (this means making the window visible). In addition, if the last active window becomes the active one after a window is closed, the former window can be made visible by SunWindow.

Again, to offer you a example of the SunWindow’s thumbnail display option of a big public screen, here’s a small graphic:

< ----  visible area of the  ----> < -- only reachable with   -- >
           public screen                 autoscroll feature
------------------------------------------------------------------
|=== Tool-                       :                               | /\
|| | manager                     :                               | v
|| | window                      :  ==========================   | i
|| |                             :  |                        |   | s
|| |                             :  |     Workbench          |   | i
|| |                             :  |                        |   | b
|---                             :  |       Window           |   | l
|                                :  |                        |   | e
|                                :  |                        |   |
|                                :  |                        |   | a
|                       ----------  |                        |   | r
|                 Sun-  |§  ##|  |                        |   | e
|                 Window|      ##|  |                        |   | a
|                       |        |  |                        |   |
|                       |%%%     |  |                        |   |
|                       |        |  --------------------------   |
|.......................----------                               | \/
|                                                                | /\
|                                                                | r
|                                                                | e
|                                                                | s
|                                                                | t
| ==============================                                 |
| |                            |                                 | o
| |          Shell             |                                 | f
| |                            |                                 |
| |                            |                                 | s
| ------------------------------                                 | c
|                                                                | r
|                                                                | e
|                                                                | e
|                                                                | n
|                                                                | \/
------------------------------------------------------------------

In this example SunWindow opens you a small display window in the right bottom of the displayed area of the screen. Now you can control the whole screen via this small display by just clicking on a window or the background to go to the desired window or screen area. Also the windows can be moved, brought to front or made visible.

1.2 System Requirements

SunWindow uses some new features introduced by version 2 of the Amiga’s OS. Therefore of course Kickstart 2.0 or higher are required. User of Kickstart 3.0 will notice some small enhancements in the preference program of SunWindow.

If you are using the <opaque movement> feature of SunWindow, a fast processor and graphic display are recommanded.

SunWindow and the preference programm will take advantage of the ‘reqtools.library’ \begin{matrix}c \\ ◯ \end{matrix} by Nico Francois. Please read the supplied document on this library in the distribution package.

1.3 How To Install SunWindow

Before you run SunWindow for the first time, you should consider the following things:

• You need the ‘reqtools.library’ in your ‘LIBS:’ drawer (V37 or higher).

• You need a valid preference file in your ‘ENV:’ drawer. The preference file usually resides in your ‘ENVARC:’ drawer and is called ‘SunWindow.prefs’. If the preference file is placed in the ‘ENVARC:’ drawer, it’ll be automatically copied to the ‘ENV:’ drawer during booting.

• The SunWindow’s preference program should have been run, and the options should have been saved once before the first use of SunWindow. This ensures that the path of the preference program is valid and therefore the preference program is accessable via a hotkey.

• And of course you should have read the manual!

If you are familar to your Amiga, I am sure you will know how to setup your system correctly for the first use. If not, follow these steps:

• Enter a shell by double clicking the shell’s icon at your ‘Tools’ drawer on the Workbench disk which comes with your Amiga.

• Insert the SunWindow distribution disk in diskdrive df0:. (This is normally the internal drive.)

• Now type the following lines in your shell, and press RETURN after each line:

       cd DF0:
       copy SunWindow#? SYS:Utilities
       copy SWPrefs#? SYS:Prefs
       copy LIBS/reqtools.library LIBS:
       copy ENV/#? ENVARC:

The first line switches to the disk drive. The following lines copy the files you need to run SunWindow. Therefore these lines will copy SunWindow to the ‘Utilities’ drawer of your SYS: drive (usually the disk with which you booted, or the hardisk partition which includes the Workbench). You may want to alter this line to place SunWindow into the ‘WBStartup’ drawer, to have SunWindow available each time you startup your system. Just replace ‘Utilities’ with ‘WBStartup’. Also the preference program ‘SWPrefs’ is copied to the ‘Prefs’ drawer of your SYS: drive. Of course this path may also be changed, to your prefered destination drawer. Be sure to copy the ENV: drawer. If you won’t do so, you might have to do a lot of initializations in the preference program.

• If you are using Kickstart 3.0 or higher, you perhapes like the only documentation available. If so, just add the following line: copy language/SunWindow.guide HELP: language where language is a replacement for the language you prefer actually only ‘English’ and ‘Deutsch’ are available). Please read the chapter about Multiview in your system documentation for further informations.

• Now close the shell, it’s no longer needed.

• Enter your ‘Prefs’ drawer, or the drawer where you copied the preference program, and start the preference program of SunWindow, called ‘SWPrefs’.

• Apply some changes to the options you want to alter or just do nothing. Leave the program by pressing the SAVE button. This saves the options in both, the ‘ENVARC:’ and ‘ENV:’ drawer. Also it ensures, that the path for the preference program is setup correctly and usable for a quick startup of the preference program via a hotkey.

• Now you are ready to run SunWindow. Just double click on its icon in the drawer where you have copied it. (Usually the ‘Utilities’ drawer, as shown above, or the ‘WBStartup’ drawer.)

2 The Features And Usage Of SunWindow

This chapter covers the usage and features of SunWindow. First, the usage is discussed and then all the features are explained. Because there are so many different features, the chapter is devided into several sections. The options are all set by the preference program ‘SWPrefs’. Read the chapter about ‘SWPrefs’, to learn more about its usage.

2.1 The Usage Of SunWindow

SunWindow is very easy to use. To start SunWindow, just do a double click on its icon. If you are a Shell/CLI fan you can also start SunWindow via a Shell/CLI. Note that SunWindow won’t start a background task, if started from a Shell/CLI. All general tooltype values are accepted as well as by a start from Workbench or CLI. Read your system manual for more information about general tooltypes. Normally no tooltype values are required, because every option is set by the preference program. Anyway, tooltype values have a higher priority over the values set by SWPrefs. If you want to run SunWindow at system startup, please place SunWindow in your ‘WBStartup’ drawer and add a DONOTWAIT tooltype to the tooltypes of SunWindow’s icon.

There are several ways of quitting SunWindow. The easiest way is to press the middle mousebutton of a three button mouse, while the thumbnail display is the active window. Of course not many user own a three button mouse and therefore there are three other possibilities of quitting SunWindow. First you can use the ‘Exchange’ program supplied by Commodore to quit SunWindow, second, you can send a CTRL-C signal, by pressing CTRL-C, if you started SunWindow via a Shell/CLI and third, you can rerun SunWindow. All quitting methods will popup a requester where you should confirm.

Mostly you don’t want SunWindow to quit, but just stop working. Use the hotkey for disabling SunWindow or the previously named ‘Exchange’ program. Because SunWindow is a commodity, all of the standard commodity functions can be controlled with ‘ Exchange’. Another circumstance of commodities in general is that if you close its window, the commodity is not quitted, instead it’s only disabled. Read your system documentation, to learn more about the internal functionality of commodities.

All other functions of SunWindow can be controlled by hotkeys or the preference program of SunWindow: SWPrefs. Go on reading the next sections, to learn more about these functions.

2.2 SunWindow and windows

With SunWindow you get a greater control over windows. The following options can be set in the preference program, to get the desired behaviour of SunWindow. All of the window options are located in the screen actions group of the preference program, because of their connection to a plain usage of SunWindow (this means, running SunWindow, without the thumbnail display).

• <center activated> This option tells SunWindow to make the active window visible on a screen by centering the display around this window. Therefore every time a window becomes active, it will be automatically made visible.

• <center opened> This option tells SunWindow to center a newly opened window. Every time a window opens on the screen, which SunWindow is just controlling (normally this is the active/frontmost screen), it will be centered.

• <center closed> This option tells SunWindow, that if a window is closed and the last active window becomes active, it will be centered. Of course setting this option is needless if you have just selected the <center activated> option.

• <active to front> Setting this option causes every active window to become the topmost window.

2.3 Screen features

SunWindow only controls public screens. This means, that you can’t use any feature of Sun-Window on a non public screen. This is not because it might be a programming problem, but it is a problem of other software integraty. If a program opens a non public screen, it just doesn’t want any other software’s windows or other software’s control on its screen. Therefore SunWindow was considered not to control these screens.

But how operates SunWindow with screens? Actually nearly not. Most of SunWindow’s screen control features can be reached via hotkeys, described in the second next section. There is only to mention, that SunWindow automatically jumps to and from screens to reach the topmost screen. Using SunWindow always causes the topmost screen to become the active screen, by activating the last active window of this screen! This is normally not handled by Intuition! Take care that this circumstance is different to the normal behaviour of your GUI.

2.4 The thumbnail display

As explained in the introduction, SunWindow is able to display the whole screen in a little window. This window is called the thumbnail display. In the thumbnail display all windows are displayed as they are positioned on the screen. With this thumbnail display you can nearly do everything to windows, as you would do it on the real screen. Therefore the thumbnail display acts as a kind of remote control for you.

If you want a specific window made visible, just click on the window in the thumbnail display. If you want the window moved to somewhere else, just click and move the window. These are only some examples, how you could use this thumbnail display. More complex actions can be performed by selecting one or more of the following options.

Several options may be set to the behaviour of this thumbnail display. But always keep in mind that all the actions you perform on the thumbnail display are also performed on the screen (of course) and therefore the options which you chose to work on windows will also affect the action! The options which you might choose for the display are listed in the preference program in the “Display actions” group and they are:

• <activate window> Every time you click on a window this window will become the active window with this option set.

• <window to front> A window you selected by clicking on it will automatically be brought to front.

• <center window> This might be the most useful option. Just click on a window and this window will be immediately made visible for you and centered on your monitor display.

• <opaque movement> You might want to set this option, if you have a fast processor paired with a fast graphics display! This option causes all window movements to be processed in realtime, this means that you will immediatly see, where the window is moved to.

• <handle backdrops> Normally there is no difference between backdrop windows and non backdrop windows. Backdrop windows are a special feature of the OS. If you set this option, you can’t make backdrop windows visible by clicking on them. Instead the area around the mousepointer is made visible! This is very useful by using the Workbench window as a backdrop window.

• <display titles> When this option is set the titles of the windows are also displayed in the thumbnail display. Unortunately sometimes the titles can’t be displayed because of a space problem, then simply nothing happens.

The following options and settings affect the lookout of the thumbnail display and are therefore located in the “Display settings” group of the preference program.

• <3D-Look> With this option set, the windows are displayed as buttons. Normally they are displayed as plain coloured boxes.

• <View border> Set this option if you want a dotted line being drawn in the thumbnail display, to indicate you the actual displayed clip.

• <Close gadget> With this option set, a close gadget is attached to the thumbnail display window. The close gadget doesn’t cause SunWindow to end (SunWindow is a commodity!), it just causes the display to be closed! Of course, also a dragbar is added, but you can’t move the display with it.

• <drag bar> Set this option if you want a dragbar attached to the thumbnail display. Now you can move the window anywhere you like, and it will stay at this relative position, everytime you move the screen, instead of the default position.

• <keep aspect> Normally you can specify the width and height of the thumbnail display in pixels. If you set this option, you can only specify the width of the display, the height is calculated, to give you the best impression of the real proportions of the screen, SunWindow is just controlling.

• <X-Coord> This value specifies the relative x-coordinate of the left top edge of the thumbnail display.

• <Y-Coord> This value specifies the relative y-coordinate of the left top edge of the thumbnail display.

• <width> This value specifies the width of the thumbnail display.

• <height> This value specifies the height of the thumbnail display.

• <Placement> Choose the default location of the thumbnail display. Setting “Individual” will take the values of <X-coord> and <Y-coord> to take affect. Setting one of the four corners will cause SunWindow to open the thumbnail 1 pixel away of the selected corner, so you can reach a possible hidden gadget. (Mostly the size, screen cycle or close gadget).

2.5 Hotkey features

Hotkeys? What is a hotkey? A hotkey is a keyboard shortcut for a specific function. This means, that if you are using a hotkey for a function you can choose, which key presses of your keyboard will act for/start this function. Therefore a hotkey is a shortcut for operations or functions which are normally only hard to get or sometimes impossible to get.

With the hotkey feature of SunWindow you get a kind of control over screens and windows you might have dreamed of. Because there are so many hotkeys available, they are again described in several subsections.

The first paragraph covers hotkeys in conjunction with window operations, the second covers hotkeys for screen operations and the third covers all those hotkeys which didn’t fit in the first two paragraphs. The hotkeys of SunWindow are implemented via the commodity system of the Amiga’s OS, so if you are not used to it, you should read the section about how to install a hotkey as well.

2.5.1 Hotkeys For Window Operations.

Below is a list of all the hotkeys of SunWindow which allow you to manipulate your windows in a way (or even better) which is normally only possible with a mouse. The first expression after an item bullet is the name of the hotkey as it would appear in the hotkey list of the preference program.

• <recenter window> Imagine this: you have just activated a Shell/CLI and want to type something in it. Now you have looked up some information in a window you can’t see, because it’s somewhere else on the screen. You only looked up this information, but didn’t activate the informative window, so you only moved the screen. O.k., now you know your information, but where is the shell? Just use this hotkey to make the active window visible again, if you moved the screen.

• <window cycle forward> With this hotkey you can cycle through all the windows of the active (topmost) screen in “forward” direction. If you selected some other features of SunWindow, this window might be made visible, brought to front or/and made activated.

• <window cycle backwards> The hotkey with the same functionality as <window cycle forward>, but of course in reverse direction.

• <enlarge window...> There are eight hotkeys for enlarging a window. With the hotkeys of SunWindow a window can be enlarged in every direction. The direction is always written in brackets. For example the hotkey for enlarging a window to its left bottom is labeled <enlarge window (left bottom)> in the hotkeys list.

• <reduce window...> Again eight hotkeys for the reverse action of enlarging a window. See hotkey description for: <enlarge window...>.

• <move window...> And again eight hotkeys for just moving a window around the screen. See also the hotkey description for: <enlarge window...>

• <toggle zip state of window> This hotkey has the same functionality as pressing the zip gadget of a window. Of course, if there is no zip gadget, you can’t toggle the zip state of a window.

• <maximize window> By pressing this hotkey you enlarge the window to its maximum dimensions. Notice, that this dimension is controlled by the program, which opened the window, and not by SunWindow.

• <minimize window> This is the reverse action to <maximize window>.

• <window to front> Just press this hotkey to bring a hidden, active window to front.

• <window to back> Again a reverse action. See <window to front>.

• <close window> This hotkey has the same affect as pressing the close gadget of a window. Of course: if there is no close gadget, you can’t close the window! Notice that a window is only closed, if the program controlling the window allows the window to be closed!

2.5.2 Hotkeys For Screen Operations.

Below is a list of all the hotkeys of SunWindow which allow you to manipulate your screen in a way (or even better) which is normally only possible with a mouse. The first expression after an item bullet is the name of the hotkey as it would appear in the hotkey list of the preference program.

• <move screen...> There are eight hotkeys for moving a screen. With the hotkeys of SunWindow a screen can be moved in every direction. The direction is always followed after the <move screen> keyword. For example the hotkey for moving a screen left down is labeled <move screen left down> in the hotkeys list.

• <cycle screens backward> With this hotkey you can switch a screen to the back. The new topmost screen will become the active screen by activating its last active window.

• <cycle screens forward> The reverse action to <cycle screens backward>.

2.5.3 Hotkeys For Miscellaneous Operations.

Below is a list of all the hotkeys of SunWindow which offer you a greater power for your system — or are just standard hotkeys. The first expression after an item bullet is the name of the hotkey as it would appear in the hotkey list of the preference program.

• <call preferences> This hotkey allows you to start the preference program of SunWindow with a simple keypress.

• <show/close view> This hotkey opens or closes the thumbnail display of the screen. This is useful if you only want to use the hotkey features, or some other features of SunWindow, which don’t rely on this display.

• <enable/disbale> This hotkey enables or disables all of SunWindow’s functions. This is always needed, if you want to change the screen mode of a screen! It is not enough, to only close the thumbnail display!

• <remember position xx> There are ten hotkeys for remembering a specific position of the mouse on a screen with the active window! The xx indicates the number (1-10) of the position you want to remember.

• <jump to position xx> These are the ten hotkeys for jumping to a previously remembered position with the <remember position xx> hotkeys. The xx indicates a number (1-10). The remembered screen is brought to front, the mouse is repositioned to the remembered position and the remember window is activated. If the screen is gone, nothing happens, if the window is gone, only the mouse is repositioned.

2.5.4 Installation Of Hotkeys.

Hokeys are installed very easily with the preference program of SunWindow. Just select the hotkey which you want to install and specify a keyboard shortcut for it. After exiting the preference program, the hotkey should be ready for usage. If the keyboard shortcut definition is wrong, a requester will popup, where you can correct the hotkey or simply alter its definition. But this requester is just a short solution! Don’t forget to correct/alter your definition in the preference program, too!

Note: a requester will also popup with a correct definition, if another hotkey with the same definition is already in use!

But how is the syntax for defining a hotkey? It’s very easy and intuitive:

• first decide, wether you want a hotkey connected with a keyboard shortcut or with the mouse. If you choose the keyboard go on reading the next item, else write done “rawmouse” as the first item in the definition string.

• now chose one or more qualifiers for your shortcut. A qualifier is one of the following items: − lshift for the left shift key on the keyboard. − rshift for the right shift key on the keyboard. − capslock for the “Caps Lock” key. − control for the control key labeled “Ctrl” − lalt for the left “Alt” key on the keyboard. − ralt for the right “Alt” key on the keyboard. − lcommand for the amiga key labeled “A”, left to the space bar. − rcommand for the amiga key labeled “A”, right to the space bar. − numericpad for a key on the numericpad (if accessable).

− repeat for key repeat (if you hold down a key long enough). − lbutton for the left mouse button. − rbutton for the right mouse button. − midbutton for the middle mouse button, if available. You can also just specify a group of qualifiers: − shift for any shift key. − alt for any “Alt” key. − command for any Amiga key (next to the space bar). Now write down this qualifier(s) in your hotkey definition string. Add a minus before those qualifiers you want to be ignored when pressed.

• Now choose wether you want upstrokes or not or both. Just ad “upstroke” to your definition string, if you want only upstrokes. Leave it, if you want only downstrokes. Add it with a preceding minus if you want both. Adding “upstroke” is optional.

• Add a highmap keycode, or an ANSI-Code to your definition string. This is also optional. A highmap keycode is one of the following items: − space for the space bar. − backspace for the backspace key. (“<-”) − f1 ... f10 for one of the “F-Keys”. − esc for the key labeled “Esc”. − tab for the tabulator key. (above the “Ctrl” key). − enter for the key labeled “Enter”. − return for the return key. − del for the key labeled “Del”. − help for the key labeled “Help”. − up, down, left, right for one of the direction keys. An ANSI-Code is just the character you get by pressing a normal key. (e.g. “a”).

Now you did everything which is needed, for a complete hotkey definition. Here some examples, which might explain it a little bit better:

• lshift alt f2: press the left shift key with any “Alt” key and the “F2” key simultaniously.

• -shift control a: press the control key and the key which produces the “a” simultaniously. If you want you can also press any shift key.

Notice: There is a bug in Kickstart 2, which prevents you using the mousebuttons. In version 3 of the Kickstart, this bug is fixed.

2.6 Miscellaneous Features

There are some features left, which didn’t fit into one of the previous sections. They are therefore placed in the “Miscellaneous” group of the SunWindow preference program (except for the <SunMouse> options, which are placed in the “Screen actions” group, because they take only affect screens.)

The features are:

• <startup window> This option causes the preference program to be run at startup of SunWindow, where you can do quick changes.

• <T:> This is a shortcut for the task priority of SunWindow. Normally different programs are run at task priority level 0. Workbench runs at level 1 (so it get’s more CPU-time in very busy situations). Background tasks may want to run at level -1, which causes them to loose the CPU in busy situations. Choose the priority you prefer.

• <C:> This is a shortcut for the commodity priority. Commodities use the same information data stream. Usually the commodity with the highest priority will get the information first. A commodity with low priority later. This may cause some lost of informations, because commodities are allowed to strip informations from the information stream. The default priority should be 0.

• <U:> This is a shortcut for the update time of SunWindow in 1/10 of a second. This will only affect the time, how fast SunWindow will react to changes on the screen, and this value may be strongly influenced by the task priority of SunWindow. A value of 3 is a nice try to start.

• <SunMouse feature> This option enables the SunMouse feature of SunWindow, if set. Sun-Mouse feature means, that always the window under the mouse will be activated.

• <SunMouse delayed> If this option is set, the window will only be activated, if the mouse stopped over a window, for a specified amount of time. You may only set this option if you have already enabled the <SunMouse feature> option.

• <SunMouse reaction...> This is the time value for the <SunMouse delayed> option in 1/10 of a second. A value of 4 is a good delay time.

3 The Preference Program: SWPrefs

When you correctly installed SunWindow, you’ll find a program called SWPrefs in the Prefs drawer of your startup disk/hardisk partition. This is the preference program for SunWindow. With this program you can set all the options SunWindow offers you. When you start SWPrefs you find a lot of options divided into four groups, concerning their functionality. There is a Display actions, a Screen actions, a Display settings and a Miscellaneous group (from top left to right bottom), additionally to the most right is a control panel attached. Each group is now explained:

The Display actions group contains all those options which concern the thumbnail display of SunWindow. The different options are descriped in the Usage chapter, section Display.

The Screen actions group contains all those options which concern the screen directly, this means that this options can be used without the thumbnail display. The different options are described in the chapter Usage, section Window, Screen and Miscellaneous.

The Display settings group contains, of course, the settings for the thumbnail display. The different settings are described in the Usage chapter, section Display.

The Miscellaneous group contains some settings which contain the internal functionality of SunWindow, here you can set priorities, etc. There is also a button labeled Keys... which takes you to the hotkey definition window. The different options in this group are described in the Usage chapter, section Miscellaneous. The hotkey definition window is described below.

The control panel on the right of the display has seven buttons. They are now explained from top to down:

• About: This button opens a little window with information of this version of SunWindow. The version number, revision number and registered user is displayed here, as well as a copyright notice.

• Load: This button reloads the last saved settings from the ENVARC: drawer.

• Load as ...: This button let’s you choose a differnt settings file, previously saved with the Save as ... button.

• Save: This button saves the settings in the ENVARC: drawer as well as in the ENV: drawer. SunWindow will get notified about the new settings. SWPrefs ends after saveing the settings.

• Save as ...: This button let’s you save the settings to a different place of your choice. SunWindow won’t get notified about this settings, until they are placed in the ENV: drawer. SWPrefs ends after saveing the settings.

• OK: This button saves the settings only in the ENV: drawer. This means, no permanent changes are done to the settings file saved on disk in the ENVARC: drawer. SWPrefs ends after saving the settings.

• CANCEL: This button quits SWPrefs and leaves all settings untouched.

Notice: Closing the preference’s window has the same effect as pressing the OK button. It saves the settings in the ENV: drawer and quit’s SWPrefs.

By pressing the Keys... button in the Miscellaneous group, another window opens with a big list view, containing all the hotkeys SunWindow offers you. Below this list view is a string gadget, where the actual defintion of a hotkey is displayed and can be edited. Two buttons labeled OK and Cancel are attached to the bottom of the window. If you want to change a hotkey just select it from the list view, by either using the mouse or the H key (unshifted scrolls down, shifted scrolls up). The definition is then displayed in the string gadget. Activate the string gadget with the mouse or by pressing E and edit the defintion. Press RETURN when finished. Go on with all the other hotkeys you want to change. When finished, press the OK button to overtake the new definitions (or just close the window), press CANCEL if you think you were doing something wrong.

Notice: The hotkeys’ definitions are not checked, wether they are correct or not. Only if you save the settings, SunWindow will check them, and popup a requester for the wrong definitions. Please correct them in the requester and with SWPrefs!

Notice: If you want to disable a hotkey, just enter nothing, or clear the actual definition.

4 Other Things You Want To Know

This chapter is devided into several sections with additional informations. The first section covers information about the registration for SunWindow, followed by sections about how to recieve updates and to get support. A special section for thanks to those people who helped me with this program is also included. The most interesting section might be Questions and Answers where commonly problems and their solutions are discussed.

4.1 How to register

SunWindow was now developed over one year during my study on University. Because I had to invest a lot of time, and can’t just work for fun, SunWindow was decided to be SHAREWARE. This means, that you can test SunWindow for a period of time (I understand up to four weeks) and if you like it you are recommanded to register, to get access to the fully working version.

A registered user will receive a keyfile. This keyfile is placed in the DEVS: drawer. With this keyfile you can use this version of SunWindow and any other version/revisions becoming available.

To register for SunWindow just fill in the following registration form, and send it to the supplied address. Your registration will be handled within a week, outside Germany registrations may take a longer time.

                        REGISTRATION FORM
I want to register for SunWindow. I am currently usin Version ____
Revision _____. I received this version by/from _________________
The registration fee is US$ 20 (outside Germany) or DM 20 (inside
Germany).
My address:
__________________________________________________________________
__________________________________________________________________
__________________________________________________________________
__________________________________________________________________
EMail: ___________________________________________________________
( ) Please register me as a group. Ich want to use additional ___
    copies of SunWindow in my group. I will therefore receive these
    copies for the special group price of 5 US$ / 5DM for each
    additional copy (*)
( ) I already registered. Please send me the latest version
    of SunWindow. I will wait up to ___weeks, if a new version
    or revision is just under development. The sharing of expenses
    is US$ 5 (outside Germany) / DM 5 (Germany only). There is
    NO fee for updates sent to you via e-mail.
( ) I want to receive my copy as fast as possible. Please send it
    to me by e-mail. (about 290kB. lha-File uuencoded)
This mail includes US$/DM _______in the form of
( ) plain money    ( ) EuroCheck
         *** SPECIAL COUNTRY CHECKS ARE NOT ACCEPTED! ***
I am using an Amiga: ___________Processor: 680__0
Kickstart:  ( )1.2  ( )1.3  ( )2.0  ( )3.0  ( ) ___
RAM: ___MB Fast, ___MB Chip
graphic card: ____________________________________________________
additional special hardware: _____________________________________
Additional informations, comments and enhancement ideas:
__________________________________________________________________
__________________________________________________________________
__________________________________________________________________
__________________________________________________________________
__________________________________________________________________
__________________________________________________________________
(*) please add the adress/name of the group.
Date: _______________Signature: ________________________________

Please send this registration form to:

        Bernhard Scholz
        Steinbruchstr. 9
        82398 Polling
        Germany

4.2 How To Receive Updates

As a registered user you will recieve a copy of the latest version of SunWindow, including a keyfile. With this keyfile you get independent from direct sendings of the author, instead you can use the actual distributed version on the aminet, or some PD-series.

If you don’t have access to the aminet or PD-series, of course you will get updates directly from the author on request. For the first time an update is available, every registered user will get an information about it. If you don’t take advantage of this new version, no further informations are sent. Of course every new version will be released on several PD-Series, and most revisions will be released in electronic form on the aminet, therefore users will get notified about the further development of SunWindow.

If you want to get an update directly from the author, there are two possible ways:

• Write to the author and include 5 US$ if you are living outside Germany, or DM 5 if you are living in Germany. The latest version will be sent to you as fast as possible.

• Wirte an electronic mail to the author, if you want to get the latest update via e-mail (lha-file uuencoded). This is the fastest way to receive an update.

The addresses for reaching the author are:

     Bernhard Scholz                 or: (e-mail)
     Steinbruchstr. 9
     82398 Polling scholz@informatik.tu-muenchen.de
     Germany

4.3 The Support You Get.

There is simply NO support for unregistered users, except that they will get notified of new version by looking on several PD-series or the aminet, because every major enhanced release will be distributed there.

Registered users may send questions, comments and wishes to the author, and they are surely considered for the next release and their questions will be answered. Please note that only those questions can be answered, by normal mail, if there is an amount of money included in your letter, to send you the answer back! E-mail will of course be the most convinient way to contact the author and to get answer.

4.4 Thanks To The People Who Helped.

During the development of SunWindow several problems occured, and I needed people for testing my software, or just commenting it. These people helped me most:

• Martin Korndörfer. For nearly every programming problem, he knowed a trick or solution. Also his “MagicMenu” is a great product! Take a look on it!

• Thomas Feistle. For correcting the documentation, and his enhancments to it. Really much work, which I hate.

• Antje Karrasch. For re-reading the docs.

• Christian Rattei. For testing SunWindow on his A3000 and finding always a bug, when I just thought every bug was fixed. He is also the guy, who makes testing with the EGS System possible. Also take a look at his fantastic disk utility “FileFairy” which will soon be released!

• Christian Scholz. He always had problems with SunWindow. This is useful for developing even better versions. He is also the guy, who painted the 8 colored icons.

• Bernhard Möllemann. For his testing, and bugreports on the good old original Amiga A1000.

• Henning Schmiedehausen. He sent most comments to the first version of SunWindow and tested it intensively. He also sent most Enforcer and Mungwall hits, just when I thought there are no more hits!

• Manfred Fichtl. For testing SunWindow on his A4000 and giving me some ideas, how to enhance it.

• Peter Kunath and Frank Riffel. Thanks also for testing and commenting. You should also register for their great work, the DeliTracker, the best sound module player ever written for the Amiga.

• Matthias Scheler. For managing #amigager on the IRC system.

4.5 The Future Of SunWindow

With your registration you’ll make a further development of SunWindow possible. SunWindow in the present state fits my personal needs and will therefore be only slightly enhanced. I will only go on developing SunWindow, if there are enough registrations and enhancement ideas. I thank all the people who will (and have already) register(ed) and support(ed) this project. Remember: SHAREWARE products depend on your support!

Some ideas for the future, which might be included in the next revisions:

• Window iconifying.

• Screen icons in a seperate window, so screens can be selected fast.

• special mousepointers (animated) for windows which don’t have their own special mousepointer design.

• Annotator tool, for different actions.

• sticky windows option

4.6 Questions and Answers

Because SunWindow is a tool, which offers you a new feeling for you computer, not everything might be common to the usual behaviour. But if you have worked a while with SunWindow and got used to it’s features, I’m sure you won’t want to miss them again! Also due to the programming difficulties with the system of the Amiga, some problems might occour. These Problem are now discussed:

• I just opened a new public screen but SunWindow’s thumbnail display still runs on the old one. Try to close the thumbnail display with the hotkey and reopen it. If the thumbnail display is still on the other screen, then your public screen is private!

• I want to change the ScreenMode, but something is still locking the screen!

SunWindow has a savety lock on the public screen it is controlling, therefore you have to disable SunWindow first. Only closing the thumbnail display isn’t enough.

• If the thumbnail display is in front of another window, I just can’t click this window to back! This is normal, due to the internal operations of Intuition. Use the hotkey for bringing a window to back instead, or close the thumbnail display.

• In conjunction with MagicMenu the systems seems to hang up sometimes.

There is no real solution, yet. But the system doesn’t hang up. Intuition just locks the menu system, after a view seconds the system is usable again.

• I set the update time to 0, but sometimes SunWindow needs a long time reacting to changes on the screen. You are running another task with a high task priority or a task which needs a lot of CPUpower. Try setting the task priority of SunWindow to 1, or simply wait until the busy task has ended.

• I am using a patch, which causes all requesters to become ReqTools requesters, but now all the requesters are automatically centered, even if I didn’t specify this! You have set the option for centering windows. Because all requesters are based on windows, this is a normal circumstance.

• I am using Directory Opus, but if I want to quit, Directory Opus pops up a requester telling to close all foreign windows?! Directory Opus belongs to those kind of programs, which don’t count on programming guidlines in some circumstances. Disable SunWindow, or put it to sleep, to get Directory Opus ending.

• I want to change the screen mode, but sometimes the system crashes!

Take care on specifying the new screen mode dimensions; it might be possible that SunWindow can’t open it’s display with this new (extrem) values!

• I am using a graphics card with the EGS system. Unfortunately SunWindow doesn’t work perfectly! The EGS system in it’s present state isn’t 100% systemcompatibel. Take care that also other programs can’t work together with EGS, because EGS isn’t error free! SunWindow was tested with EGS, but there might be still some problems not known to the author. Most Problems belong to the EGS system and not to SunWindow. Please tell the author your problems, so they can be fixed. (Only known problem: Running SunWindow without a valid SunWindow.prefs file will cause an window opening error; under EGS the system will simply crash your computer!)

5 The Distribution Package.

In this chapter you will find everything about the distribution package of SunWindow. Also information is included about redistribution, and distribution with other software.

5.1 What’s In The Distribution Package

In the distribution package of SunWindow V2.0 the following files are included. Please check whether you got all this files. Please inform the author if you didn’t get all of these files. The files are:

ENV (dir)
SunWindow.prefs
OtherIcons (dir)
SunWindow.info                   SWPrefs.info
English (dir)
ReadMeFirst.doc                  ReadMeFirst.doc.info
Registration.doc                 Registration.doc.info
ReqTools.doc                     ReqTools.doc.info
SunWindow.doc                    SunWindow.doc.info
SunWindow.dvi                    SunWindow.dvi.info
SunWindow.guide                  SunWindow.guide.info
Deutsch (dir)
ReadMeFirst.dok                  ReadMeFirst.dok.info
Registrierung.dok                Registrierung.dok.info
ReqTools.dok                     ReqTools.dok.info
SunWindow.dok                    SunWindow.dok.info
SunWindow.dvi                    SunWindow.dvi.info
SunWindow.guide                  SunWindow.guide.info
ReqTools (dir)
   libs13 (dir)
     reqtools.library
   libs20 (dir)
     reqtools.library
   Prefs (dir)
     ReqTools                         ReqTools.doc
     ReqTools.doc.info                ReqTools.info
   Install (dir)
        Workbench_2.1+ (dir)
          Install ReqTools
     Install ReqTools                 Installer.README
     Installer.README.info            Install_1.3
     Workbench_1.2-1.3                Workbench_1.2-1.3.info
     Workbench_2.0.info               Workbench_2.1+.info

          Install.info                     Prefs.info
     Deutsch.info                     English.info
     LISTING                          OtherIcons.info
     ReqTools.info                    SunWindow
     SunWindow.info                   SunWindow.small
     SunWindow.small.info             SWPrefs
     SWPrefs.info

The special file ‘LISTING’ contains the dumped directory with file sizes, etc. ‘SunWindow.small’ is a short version of SunWindow, without error messages and default settings!

5.2 The History Of SunWindow V2.0

Actually there is no history for SunWindow V2.0. This is the first release.

But for those who updated from SunWindow V1.16 or higher, here the major changes to Sun-Window V2.0:

• Preference program to control all the features, instead of a menu.

• Hotkeys added for window and screen control.

• Hotkeys added to remember positions and quickly jump back to them.

• SunMouse feature added.

• Thumbnail display enhanced, to display window titles as well.

• Greater control over the thumbnail display with dragbar, and closegadget.

• Automatic sizing of the thumbnail display, to show the screen with the right proportions.

• several internal routines completely rewritten.

5.3 The Distribution Of SunWindow

SunWindow may be distributed with other software under the following circumstances:

• the copyright belongs to me.

• the documentation must not be changed.

• I get a fully functional version of the software, if needed with registration.

• the distribution package of SunWindow is not changed.

• additional for commercial software, including SHAREWARE: − the producer has to register. − you have asked me for permission and got a licence, which is included in your Software

package

With this writing, Fred Fish and Martin Schulze get the written permission, to include Sun-Window in their PD-series (AmigaLibDisk and SaarAG-PD)

5.4 Warranty

With the use of this software, you accept the following disclaimer:

There is no warranty for the program, to the extent permitted by applicable law. Except when otherwise stated in writing the copyright holder and/or other parties provide the program "as is" without warranty of any kind, either expressed or implied, including, but not limited to, the implied warranties of merchantability and fitness for a particular purpose. The entire risk as to the quality and performance of the program is with you. Should the program prove defective, you assume the cost of all necessary servicing, repair or correction.

In no event unless required by applicable law or agreed to in writing will any copyright holder, or any other party who may redistribute the program as permitted above, be liable to you for damages, including any general, special, incidental or consequential damages arising out of the use or inability to use the program (including but not limited to loss of data or data being rendered inaccurate or losses sustained by you or third parties or a failure of the program to operate with any other programs), even if such holder or other party has been advised of the possibility of such damages.

Table of Contents

SunWindow Copying Conditions . . . . . . . . . . . . . . . . . . . . . . . . 1 1 Introduction To SunWindow . . . . . . . . . . . . . . . . . . . . . . . . . 2

1.1 The Concept Of SunWindow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 System Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.3 How To Install SunWindow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

2 The Features And Usage Of SunWindow . . . . . . . . . . . 7

2.1 The Usage Of SunWindow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2 SunWindow and windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.3 Screen features. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .8 2.4 The thumbnail display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.5 Hotkey features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

2.5.1 Hotkeys For Window Operations. . . . . . . . . . . . . . . . . . . . . . . . . 11 2.5.2 Hotkeys For Screen Operations. . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.5.3 Hotkeys For Miscellaneous Operations. . . . . . . . . . . . . . . . . . . . 12 2.5.4 Installation Of Hotkeys.. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .13

2.6 Miscellaneous Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

3 The Preference Program: SWPrefs . . . . . . . . . . . . . . . . 16 4 Other Things You Want To Know . . . . . . . . . . . . . . . . . . 18

4.1 How to register . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4.2 How To Receive Updates. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .20 4.3 The Support You Get. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.4 Thanks To The People Who Helped. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.5 The Future Of SunWindow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.6 Questions and Answers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

5 The Distribution Package.. . . . . . . . . . . . . . . . . . . . . . . . . . .24

5.1 What’s In The Distribution Package. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .24 5.2 The History Of SunWindow V2.0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 5.3 The Distribution Of SunWindow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 5.4 Warranty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26